### v2

# Variable aggregation for nonlinear optimization problems

Abstract
Variable aggregation has been largely studied as an important pre-solve algorithm for optimization of linear and mixed-integer programs. Although some nonlinear solvers and algebraic modeling languages implement variable aggregation as a pre-solve reduction, the impact it can have on constrained nonlinear programs is unexplored. In this work, we formalize variable aggregation as a pre-solve algorithm to develop reduced-space formulations of nonlinear programs. A novel approximate maximum variable aggregation strategy is developed to aggregate as many variables as possible. Furthermore, aggregation strategies that preserve the problem structure are compared against approximate maximum aggregation. Our results show that variable aggregation can generally help to improve the convergence reliability of nonlinear programs. It can also help in reducing total solve time. However, Hessian evaluation can become a bottleneck if aggregation significantly increases the number of variables appearing nonlinearly in many constraints.
Keywords:
Nonlinear optimization Presolve Graph theory
00footnotetext: We gratefully acknowledge funding from the Los Alamos National Laboratory Laboratory-Directed Research & Development program through the Center for Nonlinear Studies. Approved for unlimited release. LA-UR 25-21592.
1 Introduction
Pre-solve reductions have been important to develop linear and mixed-integer optimization solvers into generic tools that can solve real-world problems in practical settings bixby2007 . One such pre-solve reduction is to substitute an expression for a variable that is defined by an equality constraint. For example, if an optimization problem contains the constraints
$\begin{array}[]{c}x^{2}+y^{2}\leq 1\\ y=2z+1\\ \end{array}$
then the expression $12z+$ 1 may be substituted for $y$ in the first constraint to reduce the number of variables and constraints in the optimization problem. We refer to this substitution as variable aggregation and the resulting optimization problem as a reduced-space formulation. In a summary of Gurobi’s gurobimanual pre-solve methods for mixed-integer linear programming, Achterberg et al. achterberg2020 show that variable aggregation can be a source of significant speed-up, but also mention that it can be a source of numerical errors if not implemented carefully. In the context of nonlinear optimization, Bongartz bongartzthesis tests manually-chosen variable aggregations and remarks that “the reduced-space formulations reduced computational times and increased success rates of local solvers” with the caveat that “the smallest possible formulation did not necessarily result in the best computational performance” (see Sections 3.4.3.3 and 3.5 of bongartzthesis ).
For nonlinear local optimization, a generalization of variable aggregation is to remove a subset of variables and equality constraints that have a nonsingular Jacobian using the implicit function theorem. Parker et al. parker2022 and Bugosen et al. bugosen2024gibbs show that implicit function reformulations can improve reliability of interior point methods for nonlinear optimization, although these implicit function reformulations must be chosen by an expert user. While it is not easy to automatically choose general subsets of equations and variables that always have nonsingular Jacobian matrices, it is relatively straightforward to identify variables and equations that can be eliminated by explicit variable aggregation. Here, we propose automatic variable aggregation in nonlinear optimization models as a method to improve convergence reliability of an interior point solver.
Several existing software packages implement automatic variable aggregation for nonlinear optimization models. The AMPL algebraic modeling language ampl scans constraints to identify those written in the form “$y=f(x,\dots)$”. If $y$ is unbounded and has not already been aggregated, the expression $f(x,\dots)$ will be substituted for $y$ everywhere it appears. (This behavior is controlled by the substout option. See fourer1994 and Chapters 14 and 18.2 of ampl for details.) The AIMMS modeling language aimmsmanual and Pyomo modeling environment pyomobook ; pyomopaper also implement presolve options for aggregating variables, but only use equality constraints with at most two variables for these aggregations. Variable aggregation may also be performed within nonlinear optimization solvers. The Knitro knitro and CONOPT conopt solvers implement options for variable aggregation knitromanual ; conoptgamsmanual , but it is not clear what criteria these commercial solvers use to select variables and constraints to eliminate.
Despite the many variable aggregation implementations present in open-source and commercial software for nonlinear optimization, the impacts of these pre-solve strategies are not well-documented. We fill this gap by analyzing several aggregation methods inspired by those implemented in AMPL, AIMMS, and Pyomo. We analyze structural properties, solve time, and convergence reliability of optimization models before and after aggregation. Specifically, our contributions are:
  1. 1.
    We formalize variable aggregation for nonlinear optimization problems and present an easily checkable sufficient condition for a subset of variables and equality constraints to admit a valid explicit aggregation.
  2. 2.
    We present a new approximate-maximum algorithm for variable aggregation.
  3. 3.
    We demonstrate that conservative variable aggregation approaches lead to modest improvements in solve time, while aggressive variable aggregation methods increase solve time by creating expensive-to-evaluate Hessian matrices.
  4. 4.
    We demonstrate the convergence reliability benefits of variable aggregation on several parameterized nonlinear optimization test problems by sampling 121 parameter values per problem and attempting to solve each instance. For each aggregation method, we attempt to solve each problem instance. Six different aggregation methods (and the original problem) are compared by the number of instances solved for each test problem.
Improved convergence reliability of nonlinear optimization problems when aggregating variables is significant as even state-of-the-art nonlinear optimization solvers are not perfectly reliable. That is, convergence (within a pre-specified time or iteration limit) can be sensitive to initialization, scaling factors, and parameter values. Despite mature convergence theory for globalized optimization algorithms (e.g., waechter2005global ) the reasons for algorithms’ sensitivities to these values is not well-understood. By documenting the effect of variable aggregation on convergence reliability, we build an empirical foundation on which this phenomenon can begin to be understood.
The goal of this work is not to advocate for any particular aggregation method or implementation but to present a variety of approaches and analyze structure, runtime, and reliability that result when they are applied to nonlinear optimization problems.
2 Mathematical background
2.1 Nonlinear optimization
We address nonlinear optimization problems in the form given by Equation 1, where $z$ is a variable vector in $\mathbb{R}^{n}$ and $f$, $g$, and $h$ are functions with outputs in $\mathbb{R}$, $\mathbb{R}^{m_{g}}$, and $\mathbb{R}^{m_{h}}$.
$\begin{array}[]{cl}\displaystyle\min_{z}&f(z)\\ \text{s.t.}&g(z)=0\\ &h(z)\leq 0\\ &z^{l}\leq z\leq z^{u}\\ \end{array}$ (1)
Constant vectors $z^{l}$ and $z^{u}$ may have coordinates that are $\infty$- ∞ or +$\infty$+ ∞, respectively, if $z$ is unbounded in some coordinates. Lagrange multipliers for equality, inequality, lower bound, and upper bound constraints are denoted $\lambda$, $\nu$, $\gamma^{l}$, and $\gamma^{u}$. For instance, interior point methods are efficient algorithms for arriving at local solutions of problems in this form biegler2010 . They operate by iteratively solving the linear system given in Equation 2 to compute search directions $(dz,ds,d\lambda,d\nu)$.
$\left[\begin{array}[]{cccc}W+\Sigma&&\nabla g&\nabla h\\ &S^{-1}\Gamma^{s}&&I\\ \nabla g^{T}&&&\\ \nabla h^{T}&I&&\\ \end{array}\right]\left(\begin{array}[]{c}dz\\ ds\\ d\lambda\\ d\nu\\ \end{array}\right)=-\left(\begin{array}[]{c}\nabla f+\nabla g^{T}\lambda+\nabla h^{T}\nu+\left(Z_{u}^{-1}-Z_{l}^{-1}\right)\mu\mathbbm{1}\\ \nu-S^{-1}\mu\mathbbm{1}\\ g\\ h+s\\ \end{array}\right)$ (2)
Here, $W$ is the Hessian of the Lagrangian with respect to $z$ and $\Sigma=Z_{u}^{-1}\Gamma^{u}+Z_{l}^{-1}\Gamma^{l}$. Diagonal matrix $S$ has a diagonal of slack variables $s$, which are introduced along with the bound constraint $s\geq 0$ and multiplier $\gamma^{s}$ to reformulate inequality constraints. $Z_{u}$ and $Z_{l}$ are diagonal matrices of $(z^{u}-z)$ and $(z-z^{l})$, $\Gamma^{s}$, $\Gamma^{u}$, and $\Gamma^{l}$ are diagonal matrices of the corresponding Lagrange multipliers, and $\mu$ a barrier parameter chosen by the algorithm. The matrix on the left-hand-side of this equation is referred to as the Karush-Kuhn-Tucker, or KKT, matrix. Our convention for Jacobian matrices, e.g., $\nabla g$, is that rows correspond to variables while columns correspond to constraints. We note that the transpose of the Jacobian, e.g., $\nabla g^{T}$, has the same sparsity pattern as the incidence matrices we will describe in Section 2.3.
Constructing the linear system in Equation 2 requires the Hessian of the Lagrangian, the Jacobian of constraints, and the gradient of the objective function. If algebraic expression graphs for constraint and objective functions are available, these derivatives are computed by automatic differentiation (AD) of these functions griewank2008 . For most problems, the computational bottleneck of solving Equation 1 with interior point methods is factorization of the KKT matrix. However, for some problems with complicated algebraic expressions, evaluation of the Hessian of the Lagrangian, $W$, is the bottleneck.
2.2 Algebraic variables and constraints
Section 2.1 considers vector-valued functions of a variable vector $z$. In algebraic modeling language interfaces to nonlinear optimization solvers, however, variables are represented as scalar quantities that participate in algebraic expressions. These algebraic expressions define each coordinate of the functions in Equation 1. An algebraic expression can be represented programmatically as a tree where leaf nodes are (scalar) variables or constants and all other nodes are algebraic operators such as addition ($+$), subtraction (--), or multiplication (*). For example, the algebraic expression tree for the expression $13x^{2}-$ 1 is given in Figure (1).
$\hat{}$$x$22221111
Figure 1: Algebraic expression tree for the expression $13x^{2}-$ 1
Given an algebraic expression graph defining a function, useful information about the function can be computed efficiently by recursively visiting the nodes in the graph (Gay provides a good overview in gay2012 ). If common subexpressions are reused among multiple functions (or even within the same function), expressions may be represented as a directed acyclic graph (DAG), where subgraphs corresponding to common subexpressions are not duplicated.
For the aggregation methods described in Section 3, we require subroutines that process expression graphs and return (a) the set of variables that participate and (b) the set of variables that participate linearly. We denote these subroutines all_vars and linear_vars, each of which accept a scalar constraint and return the set of scalar variables that participate (linearly). These subroutines run in $\mathcal{O}(l)$ time, where $l$ is the number of nodes in the expression graph.
Table 1: Subroutines used by aggregation methods
Subroutine Inputs Outputs Time complexity Description
size Set Integer $\mathcal{O}(1)$ Return the number of elements in the set
all_vars Scalar constraint Set of variables $\mathcal{O}(l)$ Return the set of variables that participate in a constraint
linear_vars Scalar constraint Set of variables $\mathcal{O}(l)$ Return the set of variables that participate linearly in a constraint
bipartite_graph Set of variables, set of constraints Bipartite graph $\mathcal{O}(l(n_{v}+n_{e}))$ Return the bipartite graph of variable-constraint incidence
linear_bipartite_graph Set of variables, set of constraints Bipartite graph $\mathcal{O}(l(n_{v}+n_{e}))$ Return the bipartite graph of linear variable-constraint incidence
induced_subgraph Graph, set of edges Graph $\mathcal{O}(n_{e})$ Return the subgraph induced by a set of edges
maximum_matching Bipartite graph Set of edges $\mathcal{O}((n_{v}+n_{e})\sqrt{n_{v}})$ Return a maximum-cardinality matching
block_triangularize Bipartite graph, perfect matching Ordered set of sets of edges $\mathcal{O}(n_{v}+n_{e})$ Return the partition of the matching that defines the irreducible block-triangular form
2.3 Bipartite graph of variables and constraints
While expression graphs precisely represent the functional form of constraints, bipartite graphs of variables and constraints let us compute useful properties of sets of constraints (and the variables they contain).
We define a graph, $G=(V,E)$, as a set of vertices (nodes), $V$, and a set of edges, $E$. An edge $e=(u,v)$ is a pair of nodes. A bipartite graph, $G=(A,B,E)$, contains two disjoint sets of nodes $A$ and $B$ where, for every edge $(a,b)\in E$, $a$ is in $A$ and $b$ is in $B$. A node $v$ and edge $e$ are incident if $e$ contains $v$. The degree of a node is its number of incident edges. A subgraph induced by a subset of nodes $V_{s}\subset V$ is the graph $(V_{s},E_{s})$, where $E_{s}$ is the set of edges with both nodes in $V_{s}$. A subgraph induced by a subset of edges $E_{s}\subset E$ is the subset induced by the set of nodes $V_{s}$, where $V_{s}$ is the set of incident nodes of edges in $E_{s}$.
A bipartite graph may be represented as an incidence matrix, where rows correspond to one set of nodes and columns correspond to the other. An entry exists for the sparse matrix row-column pair if an edge exists between the corresponding nodes in the bipartite graph.
Figure 2: Bipartite graph and incidence matrix corresponding to variables and equations with a maximum matching highlighted. [^1] Figure reproduced from parker2023dulmage .
A matching, $\mathcal{M}$, is a set of edges, no two of which share a node. A maximum matching is a matching of the largest cardinality possible for a given graph, and a perfect matching is a matching that covers every node in the graph. A maximum matching in a bipartite graph may be computed in $\mathcal{O}((n_{v}+n_{e})\sqrt{n_{v}})$ time, where $n_{v}$ is the number of nodes and $n_{e}$ is the number of edges, by the algorithm of Hopcroft and Karp hopcroft1973 . Not all bipartite graphs have a perfect matching, but a subgraph induced by a matching $\mathcal{M}$ always has a perfect matching ($\mathcal{M}$ itself). A maximum matching is not unique, even if it is perfect; that is, many different matchings may have the same (maximum) cardinality.
Figure 3: Block triangular form of an incidence matrix. Here, $T_{1},\dots,T_{n_{b}}$ are the submatrices corresponding to subgraphs induced by subsets in the block triangular partition. [^2] Figure adapted from parker2023dulmage .
The incidence matrix of a bipartite graph that admits a perfect matching can be permuted into irreducible block lower triangular form by an algorithm given by Duff and Reid duff1978implementation ; duff1978algorithm . This is done by a subroutine, called block_triangularize in Table 1, that accepts a bipartite graph and a perfect matching thereof and returns an ordered partition $\mathcal{T}$ of the perfect matching. Permuting the rows and columns of the incidence matrix into the order defined by the partition will put it into irreducible block lower triangular form. Each subset of edges $T\in\mathcal{T}$ corresponds to a diagonal block of the incidence matrix. The diagonal blocks are irreducible in the sense that submatrices defined by each $T\in\mathcal{T}$ cannot be further partitioned into block triangular form. That is, if $\mathcal{T}_{T}$ is the block triangular partition of block $T$, ${\tt size}(\mathcal{T}_{T})=1$. This property of irreducibility is known as the Strong Hall Property scott2023 ; coleman1986 . An illustration a block triangular partition is given in Figure 2, and details about the block triangularization algorithm are given in Appendix A.
Given a set of scalar variables, $X$, and a set of scalar constraints, $C$, a bipartite graph may be constructed where one set of nodes is the set of variables and the other is the set of constraints. An edge exists between a variable node $x$ and a constraint node $c$ if $x$ participates in $c$. Figure 1 illustrates a bipartite graph and incidence matrix corresponding to three-variable, two-constraint system. Alternatively, the edges may be restricted to only the pairs corresponding to linear variable-constraint incidence. An edge corresponding to linear variable-constraint incidence is referred to as a linear edge, while any other edge is a nonlinear edge. Subroutines that return or accept bipartite graphs of variables and constraints are used by the aggregation methods in Section 3, and are listed in Table 1, where $n_{v}$ and $n_{e}$ are the numbers of vertices and edges in a graph.
If the bipartite graph of a set of variables and constraints admits a perfect matching, it can be used to determine whether the variables and constraints can be permuted to have a lower triangular Jacobian. Such a permutation exists if and only if the irreducible block triangular partition $\mathcal{T}$ contains no subsets with cardinalities greater than one. This is due to the irreducibility of the subsets of the block triangular partition.
2.4 Variable aggregation
Given an optimization problem in the form of Equation (1), it may be possible to partition the equality constraint functions $g=(\bar{g},\tilde{g})$ and variables $z=(u,v)$ such that the Jacobian $\nabla_{v}\tilde{g}$ is always nonsingular (over a domain of interest). In this case, by the implicit function theorem, there exists a function $v=\tilde{g}_{v}(u)$ that satisfies $\tilde{g}(u,\tilde{g}_{v}(u))=0$. The function $\tilde{g}_{v}$ is exploited in reduced space optimization methods such as parker2022 ; pacaud2022 ; pacaud2024 to eliminate variables and equality constraints (here, $v$ and $\tilde{g}(u,v)=0$) from the optimization problem. The resulting reduced-space optimization problem is given by Equation 3.
$\begin{array}[]{cl}\displaystyle\min_{u}&f(u,\tilde{g}_{v}(u))\\ \text{s.t.}&\bar{g}(u,g_{v}(u))=0\\ &h(u,g_{v}(u))\leq 0\\ &u^{l}\leq u\leq u^{u}\\ &v^{l}\leq g_{v}(u)\leq v^{u}\\ \end{array}$ (3)
A subset of variables $v$ and constraints $\tilde{g}(u,v)=0$ to eliminate is represented programmatically as a matching, or a set of variable-constraint edges in the bipartite incidence graph. We refer to a matching that is used for variable aggregation as an elimination order or an aggregation set. While we refer to Problem (3) as a reduced-space optimization problem, it may have more constraints (where “constraint” is defined to exclude variable bounds) than the original problem. This is because bounds on the eliminated variables, $v$, in the original problem are converted to inequality constraints in the reduced-space problem. While aggregating a variable eliminates an equality constraint, it may introduce up to two additional inequality constraints if the variable has upper and lower bounds.
Identifying whether a general subset of variables and constraints $v$ and $\tilde{g}(u,v)=0$ is nonsingular over a domain of interest involves solving a global optimization problem, which is beyond the scope of our intended presolve application. However, it is sufficient (but not necessary) for the constraints to have the form given by Lemma 1.
Lemma 1
Let
$\tilde{g}(u,v)=v-\tilde{g}^{\rm def}(u,v)=0$
be a subset of constraints in Equation 3, where $\nabla_{v}\tilde{g}^{\mathrm{def},T}$ is strictly lower triangular. Then $\nabla_{v}\tilde{g}^{T}$ is nonsingular everywhere and the subsets of variables and constraints $v$ and $\tilde{g}$ may be eliminated to form a reduced-space optimization problem.
Proof
$\nabla_{v}\tilde{g}^{T}$ is lower triangular with a diagonal of all ones, and therefore is nonsingular.
A subset of constraints can easily be checked for this form by inspecting algebraic expression graphs and the bipartite incidence graph. Here, $v$ is the vector of defined variables and $g^{\rm def}$ is the defining function (composed of defining expressions).
Constraints with this form have the additional advantage that a reduced-space optimization problem may be constructed explicitly by recursively substituting each eliminated variable $v_{i}$ with the corresponding coordinate of the defining function $\tilde{g}^{\rm def}_{i}$. Because of strict lower triangularity of $\tilde{g}^{\rm def}$, the eliminated variable $v_{i}$ does not appear nonlinearly in the equation used to eliminate it, so it is never necessary to solve for $v_{i}$ by an iterative (or implicit) equation solving method.
While one can easily check whether constraints have the form of Lemma 1, in a general optimization problem there are a combinatorial number of subsets of variables and constraints that could have this form. There are several approaches one can take to decide which subset to choose for aggregation. One approach, used by AMPL with the substout=1 option, is to try to eliminate as many constraints as possible without breaking strict lower triangularity of $\nabla_{v}\tilde{g}^{\mathrm{def},T}$. A more conservative approach, taken by AIMMS and Pyomo, is to only eliminate linear equality constraints with at most two variables. We refer to this as a linear-degree-2 elimination strategy.
Other approaches could be considered. For instance, one can set an upper bound on the number of variables in an eliminated constraint that is greater than two. Nonlinear constraints with only certain functional forms, such as low-order polynomials, could be considered. Additionally, the “target constraints” into which $\tilde{g}^{\rm def}_{i}$ is substituted could be inspected to avoid aggregations that convert linear constraints into nonlinear constraints (as suggested by Amarger et al. amarger1992 ). In this work, we consider aggregation strategies involving constraints with at most two variables as well as aggregation strategies which attempt to eliminate the maximum number of variables. The specific strategies we implement are covered in Section 3.
3 Aggregation methods
While aggregating the maximum number of variables may be valuable for some problems, making significant modifications to problem structure may compromise solve time or convergence reliability for others. For this reason, we also propose aggregation strategies that preserve certain properties of the Jacobian of remaining constraints $\bar{g}$ and $h$. This section details the variable aggregation methods that we compare in Section 5. Each elimination method is given as an algorithm that accepts sets of all variables $X$ and all constraints $C$ and returns a set of variables and constraints to aggregate as a matching $\mathcal{M}$ of the corresponding nodes in the bipartite incidence graph.
3.1 Approximate maximum elimination
A natural approach is to aggregate as many variables and constraints that have the form of Lemma 1 as possible. This is similar to the approach suggested by Christensen christensen1970 . However, identifying the maximum aggregation set (according to the form given by 1) is NP-complete.
Theorem 3.1
Let $X$ be a set of scalar variables, $C$ a set of scalar algebraic constraints, and $G$ the corresponding bipartite incidence graph. Let $\mathcal{M}$ be a matching of linear edges of $G$ where the incidence matrix of the subgraph induced by $\mathcal{M}$, $G_{\mathcal{M}}$, is lower triangular. We refer to the problem of identifying $\mathcal{M}$ with maximum cardinality as the maximum aggregation set problem. The maximum aggregation set problem is NP-complete.
Proof
The proof is by reduction to the minimum tearing problem. Let $X$ and $C$ be sets of variables and linear constraints with the same cardinalities that admit a perfect matching. It is sufficient to show that identifying a maximum-cardinality matching of $X$ and $C$ with a lower triangular induced submatrix is NP-complete. Identifying this maximum-cardinality matching is equivalent to identifying a the largest square submatrix of $X$ and $C$’s incidence matrix that is lower triangular, which is the tearing problem as described by Carpanzano carpanzano2000order . Carpanzano proves that this problem is NP-complete by reduction to the minimum feedback arc set problem (Problem 8 on Karp’s list of NP-complete problems karp1972reducibility ). Because this problem is a special case of our maximum aggregation set problem, the maximum aggregation set problem is NP-complete as well.
Because identifying a maximum aggregation set is NP-complete, we do not attempt to solve it exactly, but instead develop heuristics to identify large aggregation sets.
3.1.1 Greedy heuristic
The first heuristic we implement constructs an elimination order by iterating over constraints and checking whether there exists a variable that appears (a) linearly and (b) does not participate in any constraint already in the elimination order. If these two conditions are met, the variable-constraint pair is added to the elimination order. This approach is given by Algorithm 1 and is referred to as the greedy aggregation method, or GR for short. By construction, Algorithm 1 returns a set of variable-constraint pairs in the form required by Lemma 1. This method in inspired by the description of the aggregation phase in AMPL’s presolver given in Chapter 18.2 of ampl .
Algorithm 1 : greedy_aggregation_set
1:Inputs: Sets of variables $X$ and constraints $C$
${}\mathcal{M}=\{\}$
3:$S=\{\}$
4:for $c\text{~{}\bf in }C$ do
5:     for $x\text{~{}\bf in }\tt linear\_vars(c)$ do
6:         if $x\notin S$ then
7:              $\mathcal{M}\leftarrow\mathcal{M}\cup\{(x,c)\}$
8:              $S\leftarrow S\cup\text{\tt all\_vars}(c)$
9:              break
10:         end if
11:     end for
12:end for
13:Return: $\mathcal{M}$
Let set $\mathcal{M}$ be the set of variables and the corresponding constraints that are to be eliminated and, let set $S$ be the set of all variables that appear in the constraints in $\mathcal{M}$. Given a set of variables $X$ and constraints $C$, the greedy heuristic algorithm consists of the following steps:
  1. 1.
    Initialize set ${}\mathcal{M}=\{\}$ and set $\mathcal{S}=\{\}$.
  2. 2.
    For each constraint in $C$, compute a set of variables that appear linearly in the constraint using the subroutine ${\tt linear\_vars}$.
  3. 3.
    For each variable $x$ in the set of linear variables in a constraint, if $x$ is not already in $\mathcal{S}$:
    1. (a)
      Add the variable $x$ and the corresponding constraint $c$ to $\mathcal{M}$. $x$ has not been used by any previous constraint, so adding it will preserve lower triangularity.
    2. (b)
      Add all the variables appearing in constraint $c$ to set $\mathcal{S}$. This prevents any of these variables, which would need to precede $x$ in a lower triangular ordering, from being added later in the algorithm.
3.1.2 Matching-based algorithm
Figure 4: Illustration of the linear matching algorithm for the given equation system. Middle: A maximum matching of linear variable-constraint pairs ($\mathcal{M}_{L}$ in Algorithm 2). Right: A subset of the matching with a lower-triangular incidence matrix ($\mathcal{M}_{T}$ in Algorithm 2).
The second heuristic we implement decomposes the maximum-elimination problem into two well-studied subroutines. The first is to compute a maximum matching on the linear bipartite graph of variables and constraints. To generate a subset of this matching with a lower-triangular incidence matrix, we construct the subgraph of the full bipartite graph (containing nonlinear edges) induced by this matching, then compute the irreducible block triangular partition of this subgraph. In this block triangularization, any diagonal blocks of size greater than one-by-one break lower triangularity. To approximate the largest lower triangular subset of variable-constraint pairs in each of these diagonal blocks, we use the greedy method defined by Algorithm 1. We refer to this approach as the linear matching aggregation method, or LM for short.
The problem of identifying the largest lower triangular submatrix of a square, well-constrained incidence matrix is the well-studied tearing problem steward1965partitioning ; elmqvist1994tearing . Sophisticated heuristics for solving this problem are available fletcher1993ordering ; tauber2014tearing , and an exact algorithm (that is not efficient in the worst case) has been developed for the related minimum feedback arc set problem baharev2021mfasp . We use a comparatively simple heuristic and defer the integration of more sophisticated tearing methods to future work.
Figure 4 illustrates the linear-matching algorithm. Given a set of equations, a maximum matching is computed with the matched edges highlighted in red in the incidence matrix. However, all the matched edges cannot be eliminated since eliminating both $x_{1}$ and $x_{4}$ is not possible since the expression for $x_{4}$ contains $x_{1}$ and the expression for $x_{1}$ contains $x_{4}$. Therefore, the largest lower triangular subset is computed for aggregation. The largest lower triangular subset is highlighted with the orange rectangle and the red edges indicate the variable constraint edges that will be aggregated.
Pseudocode for the linear_matching method is given by Algorithm 2. Let $\mathcal{M}_{T}=\{\}$, be the set of variables and corresponding constraints to be eliminated. Given a set of equations with variables $\mathcal{X}$ and constraints $\mathcal{C}$, the matching based algorithm consists of the following steps:
  1. 1.
    Compute a bipartite graph $G$ of the set of equations consisting of all edges.
  2. 2.
    Compute a linear bipartite graph $G_{L}$ of the equations consisting of only linear edges.
  3. 3.
    Compute a maximum matching $\mathcal{M}_{L}$ on the linear bipartite graph
  4. 4.
    Compute the subgraph $G_{\mathcal{M}}$ of the bipartite graph $G$ induced by the maximum matching $\mathcal{M}_{L}$
  5. 5.
    Compute a block triangular partition $\mathcal{T}$ of the subgraph induced by the matching using the $\tt block\_triangularize$ subroutine.
  6. 6.
    For each block in $\mathcal{T}$:
    1. (a)
      If the size of the block is $1$, add the variable constraint pair in the block to $\mathcal{M}_{T}$
    2. (b)
      If size of the block is $>1$, the $\tt greedy$ heuristic (Algorithm 1) is used on the block to extract the variable constraint pairs to be added to $\mathcal{M}_{T}$.
By extracting a lower triangular submatrix from each block in the block triangular partition, the submatrix induced by the new matching $\mathcal{M}_{T}$ is lower triangular. Because the block triangular partition was computed from a matching of only linear edges, $\mathcal{M}_{L}$, the reduced matching $\mathcal{M}_{T}$ contains only linear edges as well. The matchings $\mathcal{M}_{L}$ and $\mathcal{M}_{T}$ are illustrated in Figure 4.
The runtime of this algorithm is dominated by either the runtime of computing the maximum linear matching $M_{L}$, or, if $G_{L}$ is small, the runtime of constructing $G_{L}$ in the first place.
Algorithm 2 : linear_matching_aggregation_set
1:Inputs: Sets of variables $X$ and constraints $C$
2:$G=\text{\tt bipartite\_graph}(X,C)$
3:$G_{L}=\text{\tt linear\_bipartite\_graph}(X,C)$
4:$\mathcal{M}_{L}=\text{\tt maximum\_matching}(G_{L})$
5:$G_{\mathcal{M}}=\text{\tt induced\_subgraph}(G,\mathcal{M}_{L})$
6:$\mathcal{T}=\text{\tt block\_triangularize}(G_{\mathcal{M}},\mathcal{M}_{L})$
7:$\mathcal{M}_{T}=\{\}$
8:for $T\text{~{}\bf in }\mathcal{T}$ do
9:     if $\text{\tt size}(T)=1$ then
10:         $\mathcal{M}_{T}\leftarrow\mathcal{M}_{T}\cup T$
11:     else
12:         $\mathcal{M}_{T}\leftarrow\mathcal{M}_{T}\cup\text{\tt greedy}(T)$
13:     end if
14:end for
15:Return: $\mathcal{M}_{T}$
We note that the cardinality of the maximum matching on the linear bipartite graph, $\mathcal{M}_{L}$ in Algorithm 2, is an upper bound on the number of variables that can be aggregated using constraints in the form given by Lemma 1. This algorithm therefore has the advantage that the suboptimality of its solution, relative to a true maximum-cardinality aggregation, can be bounded.
We also note that $\text{\tt greedy}(T)$ on line 12 of Algorithm 2 returns a set of at least one edge, as there are at least $\text{\tt size}(T)$ edges of linear variable-constraint incidence in the bipartite graph of variables and constraints in $T$. This follows from the fact that $\mathcal{T}$ is constructed from a matching $\mathcal{M}_{L}$ with only linear edges. This gives us $\text{\tt size}(\mathcal{T})$, or the number of diagonal blocks in the block triangular form, as a lower bound on the cardinality of the aggregation returned by this algorithm. This result is summarized in Theorem 3.2.
Theorem 3.2
Let $X$ be a set of variables and $C$ a set of constraints, with bipartite graph $G$ and linear bipartite graph $G_{L}$. Let $n_{\rm match}$ be the cardinality of $\mathcal{M}_{L}$, a maximum matching of the linear bipartite graph. Let $G_{\mathcal{M}}$ be the subgraph of $G$ induced by $\mathcal{M}_{L}$, with a block triangular partition $\mathcal{T}$ of cardinality $n_{\rm block}$. Let $\mathcal{A}$ be the matching returned by Algorithm 2 that is used for aggregation, with cardinality $n_{\rm agg}$. Then
$n_{\rm block}\leq n_{\rm agg}\leq n_{\rm match}$
Proof
Each block $T\in\mathcal{T}$ contributes at least one edge to the matching $\mathcal{A}$. If $\texttt{size}(T)=1$, it contributes exactly one edge. If $\texttt{size}(T)>1$, it contributes at least one edge as the first constraint encountered in greedy contains at least one linear variable (the variable matched with this constraint in $\mathcal{M}_{L}$). Therefore, $n_{\rm block}\leq n_{\rm agg}$.
Each block $T\in\mathcal{T}$ contributes at most ${\tt size}(T)$ edges to $\mathcal{A}$. This is by construction, as greedy contributes at most one edge per constraint provided as input, and there are ${\tt size}(T)$ constraints provided as input. If ${\tt size}(T)=1$, then exactly ${\tt size}(T)$ edges are contributed to $\mathcal{A}$. The number of edges in $\mathcal{A}$ is therefore at most $\sum_{T\in\mathcal{T}}{\tt size}(T)$, which equals $n_{\rm match}$ as $\mathcal{T}$ partitions $\mathcal{M}_{L}$. That is,
$n_{\rm agg}\leq\sum_{T\in\mathcal{T}}{\tt size}(T)=n_{\rm match}$
We note that the lower bound in Theorem 3.2 is specific to the particular linear maximum matching that was computed by the maximum_matching subroutine. Computing a maximum matching that induces a highly decomposable block triangular form has been studied by Lima et al. lima2006 using metaheuristic methods. We leave the integration of heuristic or exact methods for choosing a “highly decomposable matching” to future work.
3.2 Fixed-variable aggregation
A variable is fixed if it is the only variable in a linear equality constraint. A simple aggregation strategy is to replace fixed variables with the value defined by their single-variable constraints. Algorithm 3 describes the steps used to generate an ordered set of variables and constraints that can be eliminated. The filter_constraints_fixed_variable subroutine filters linear, degree-one constraints. That is, constraints of the form $y=a$ where $a$ is a constant pass through the filter.
We refer to this algorithm as the linear degree-1 aggregation method, or “LD1” for short. The subsystem containing the filtered constraints is sent to the linear_matching subroutine given in Algorithm 2 to generate the ordered set of variables and constraints that are eliminated. The linear_matching subroutines prevents situations where two constraints are assigned to eliminate a single variable. In this case of constraints each containing only a single variable, this would imply either that one of the constraints is redundant or that the model is infeasible.
Because aggregating fixed variables and their defining constraints may reduce the number of variables in other constraints, doing so may lead to more fixed variables in the model. For this reason, this strategy is always implemented recursively until there are no fixed variables left to eliminate. Fixed-variable aggregation reduces the number of variables per constraint and the complexity of expression graphs. It reduces the number of non-zeros in the Jacobian as it doesn’t introduce any additional variables into the problem.
Algorithm 3 : degree_1_aggregation_set
1:Inputs: Sets of variables $X$ and constraints $C$
2:$C^{\prime}=\text{\tt filter\_constraints\_fixed\_variable}(C)$
3:$\mathcal{M}=\text{\tt linear\_matching}(X,C^{\prime})$
4:Return: $\mathcal{M}$
Table 2: Subroutines used to filter constraints
Subroutine Inputs & Outputs Example acceptable constraint
filter_constraints_degree2 Set of constraints $y=2x^{2}$
filter_constraints_linear_degree2 Set of constraints $y=2x+3$
filter_constraints_equal_coefficient Set of constraints $y=x+4$
filter_constraints_fixed_variable Set of constraints $y=1$
3.3 Structure-preserving aggregation
While aggregating as many variables and constraints as possible leads to a problem with fewer variables (and often fewer constraints), the remaining constraints may have more variables on average and more complicated nonlinear algebraic expressions. To limit increases in density and complexity of the reduced-space problem, we implement methods that attempt to preserve the structure of the remaining constraints.
These methods aggregate using constraints with at most two variables, as aggregating a variable using such a constraint will never increase the number of variables in other constraints. The structure-preserving methods differ by what other properties, e.g. linearity, they require these defining constraints to have. The algorithms to identify aggregation sets rely on constraint filters that are summarized in Table 2, in order from most permissive to most restrictive. We note that more permissive filters, e.g., filter_constraints_degree2, still return “less general” constraints, such as a “variable-fixing” constraint $y=1$.
Aggregating variables using degree-2 constraints may decrease the number of variables in other constraints. For example, aggregating
$y\leftarrow x^{2}+2$
into the constraint
$w=x+y$
yields the new constraint
$w=x^{2}+x+2,$
which has only two variables as opposed to the original three. As in this example, such aggregations may lead to more two-variable constraints in the reduced-space model. The structure-preserving aggregation methods in this section are implemented using the following procedure:
  1. 1.
    Aggregate as many fixed variables as possible by recursively calling Algorithm 3 and applying the resulting aggregation.
  2. 2.
    Aggregate as many variables as possible using two-variable constraints by recursively applying a structure-preserving method.
3.3.1 Incidence-preserving aggregation
The simplest structure-preserving strategy is to aggregate as many variables as possible using constraints with at most two variables, at least one of which appears linearly. We refer to this as the degree-2 aggregation method, or D2 for short. Because each defining constraint contains two variables, this method cannot increase the number of variables in any constraint in the reduced-space model. The method is described by Algorithm 4, where the filter_constraints_degree2 subroutine allows all degree-2 constraints to pass through the filter. The linear_matching subroutine ensures that aggregated variables appear linearly in their defining constraints. While this strategy does not increase the number of variables per constraint, it may introduce nonlinearities into constraints that were previously linear.
Algorithm 4 : degree_two_aggregation_set
1:Inputs: Sets of variables $X$ and constraints $C$
2:$C^{\prime}=\text{\tt filter\_constraints\_degree2}(C)$
3:$\mathcal{M}=\text{\tt linear\_matching}(X,C^{\prime})$
4:Return: $\mathcal{M}$
3.3.2 Linearity-preserving elimination
A more conservative strategy is to aggregate variables using only linear constraints with two variables. We refer to this as the linear degree-2 aggregation method, or LD2 for short. This method does not increase the number of variables per constraint, and cannot convert a linear constraint into a nonlinear constraint. This method is defined by Algorithm 5, where filter_constraints_linear_degree2 allows linear constraints with two variables, i.e. constraints of the form $y=ax+b$ where $x$ and $y$ are scalar variables and $a$ and $b$ are constants, to pass through the filter. While the number of variables per constraint does not increase, constraints’ linear coefficients can change. For example, performing the aggregation
$y\leftarrow 100x+1$
in the constraint
$w=2y$
yields the new constraint
$w=200x+2,$
which has linear coefficients of $[1~{}-200]$ (when both variables are brought to the left-hand-side) instead of the original coefficients of $[1~{}-2]$. A similar algorithm is implemented in the Pyomo nl_v2 writer with the linear_presolve=True option.
Algorithm 5 : linear_degree_two_aggregation_set
1:Inputs: Sets of variables $X$ and constraints $C$
2:$C^{\prime}=\text{\tt filter\_constraints\_linear\_degree2}(C)$
3:$\mathcal{M}=\text{\tt linear\_matching}(X,C^{\prime})$
4:Return: $\mathcal{M}$
3.3.3 Equal coefficient degree-2 aggregation
The most restrictive structure-preserving aggregation strategy attempts to preserve entries in the Jacobian matrix of remaining constraints. To do this, we aggregate variables using linear constraints with two variables, where the variables have coefficients with the same magnitude. For example, a constraint
$y=x+a$
may be used to eliminate $y$. If $x+a$ is substituted into constraints that do not already contain $x$, Jacobian values in the reduced-space model will be unchanged. That is, the derivative of the original constraint with respect to $x$ is the same as the derivative of the new constraint with respect to $y$. If $x+a$ is substituted into constraints that already contain $x$, the new constraint will have one fewer variable than the original constraint, and the Jacobian entry corresponding to the derivative with respect to $x$ may change. Like the linear degree-2 aggregation strategy, this strategy preserves linearity.
We refer to this as the Equal coefficient degree-2 aggregation method, or ECD2 for short. This method is described by Algorithm 6. Here, the filter_constraints_equal_coefficient subroutine allows constraints of the form $y=x+a$ to pass through the filter. The subsystem formed by the filtered constraints and the variables that appear in the filtered constraints is passed to the linear_matching subroutine to obtain an ordered set of variables and constraints to eliminate.
Algorithm 6 : equal_coefficient_aggregation_set
1:Inputs: Sets of variables $X$ and constraints $C$
2:$C^{\prime}=\text{\tt filter\_constraints\_equal\_coefficient}(C)$
3:$\mathcal{M}=\text{\tt linear\_matching}(X,C^{\prime})$
4:Return: $\mathcal{M}$
4 Test problems
This section briefly describes the test problems on which we evaluate the methods described in Section 3. To study the effect of aggregation method on solve time, we solve nominal instances of each of the four test problems after applying each aggregation method. To study the affect of aggregation method on convergence reliability, for each of the three test problems that support parameters and each of the aggregation methods, we sample two parameters at eleven uniformly spaced points each (121 total points) and perform a parameter sweep.
4.1 Distillation column dynamic optimization
Distillation is a widely employed separation technique based on the difference in relative volatility. Dynamic distillation column models play an important role in optimally controlling the distillation column. The dynamic model consists of differential algebraic equations (DAEs) for mass balance that scale with the number of trays, control horizon, and the time discretization along with non-linear constraints linking the mole fractions. The resulting problem is a nonlinear dynamic optimization problem. The full formulation for this model (DIST) may be found in parker2022 .
4.2 Moving bed reactor optimal operation
A moving bed (MB) reactor is a two-phase chemical reactor in which gas and solid material streams flow in counter-current directions, reacting as they come in contact with each other. The inlet flow rates of gas and solid streams may be adjusted to achieve desired product compositions or temperatures. As a test problem, we consider a moving bed reactor with methane and iron oxide inlet streams operating at steady state conditions. The methane reduces the iron oxide, producing carbon dioxide and water vapor via the reaction
[…]
This reactor is modeled with nonlinear, one-dimensional differential-algebraic equations discretized along the length domain of the reactor, as described by Ostace et al. ostace2018 and Okoli et al. okoli2020 and implemented in the IDAES model library lee2021idaes . In this test problem, we penalize the deviation from a set-point of 95% oxygen carrier conversion and an inlet solid flow rate of 591 kg/s.
4.3 Pipeline network dynamic optimization
The gas pipeline network optimization problem (PIPE) is a constrained NLP consisting of a system of DAEs modeled in IDAES lee2021idaes . The goal is to minimize the gas transmission cost while meeting the customer demands and satisfying the mass and momentum balance equations for gas transport NAIK20231847 . IDAES provides a modular approach to model the gas pipeline flowsheet by leveraging existing unit models and material properties. However, this approach also leads to additional redundant variables and constraints that link the individual unit models and property packages to build the overall flowsheet.
4.4 AC optimal power flow
Alternating current optimal power flow (ACOPF) describes the task of selecting power generator dispatch levels and bus voltages to minimize cost while satisfying the AC power flow equations over a network cain2012 . We use as a test problem a 4,917-bus case from the PGLib-OPF test set pglib . The test problem data is parsed and converted into a Pyomo model by Egret knueven2019 .
5 Computational results
This section describes the results of applying the methods of Section 3 to the problems of Section 4 before solving with IPOPT ipopt version 3.14.17 with function and derivative evaluations performed by the AMPL Solver Library (ASL) asl . All test problems are implemented in Pyomo, and, where applicable, differential equations are discretized using Pyomo.DAE pyomodae . The methods to compute aggregation sets, described in Section 3, are implemented using Incidence Analysis parker2023dulmage , a Pyomo extension for analyzing incidence graphs. Structural results are presented in Section 5.1, runtime results are presented in Section 5.2, and convergence reliability results are presented in Section 5.3.
Structural results refer to solver-agnostic model properties that do not depend on particular values of variables and parameters at which the model instance is evaluated. Examples of structural results include the numbers of constraints and variables, and the number of nonzero entries in the KKT matrix. As structural results do not depend on the chosen parameter values, we compute these results for nominal instances of each test problem with each aggregation method applied. Structural results are computed using Incidence Analysis, PyNumero pynumero , and the Pyomo .nl writer’s AMPLRepn data structure to investigate the structure of incidence matrices, derivative matrices, and expression graphs.
Runtime results are those obtained by solving a particular instance of a model with a particular solver (here, IPOPT), such as solve times and numbers of iterations. To limit the impact of difficult instances, which may have extremely long solve times, and to avoid questions about how to penalize instances that don’t converge, we compare detailed timing statistics only for nominal instances of each test problem with each aggregation method applied. Runtime results are obtained by timing evaluation callbacks in the CyIpopt interface to IPOPT cyipopt .
Convergence reliability results are computed over a sweep of parameter values for the three test problems that support parameters (distillation, moving bed, and pipeline optimization problems). For each of these problems, two parameters in the optimization formulation are chosen, along with ranges over which we would like to vary these parameters. Parameter ranges are chosen to yield challenging optimization problems that may not converge within the 3000-iteration limit. For each parameter, we sample eleven uniformly spaced values for a total of 121 parameter combinations at which we solve the optimization problem. We perform this parameter sweep after applying every aggregation method and compare the results.
The code used to produce the results presented in this section may be obtained at https://github.com/Robbybp/variable-elimination. The results presented are collected on HPE ProLiant XL170r server nodes with two Intel 2.1 GHz CPUs and 128 GB of RAM running the RHEL8 Linux distribution and using Python 3.11.5.
5.1 Structural results
We first compare the effect of variable aggregation on model structure. We define model structure as solver-agnostic properties of the optimization model that do not depend on particular variable values, such as numbers of variables and constraints. A summary of model structures after aggregating with each method is given in Table 3.
Here, “Var.” is the number of variables in the model, “Con.” is the number of constraints, “Elim.” is the number of variables that were eliminated, “NNZ/Con.” is the average number of (structural) nonzero entries in the Jacobian matrix per constraint, “Lin. NNZ/Con.” is the average number of nonzero entries in the Jacobian matrix corresponding to linear variable-constraint incidence, and “Hess. NNZ” is the number of nonzeros in the Hessian of the Lagrangian.
Table 3: Structural properties of models after applying each aggregation method
Model Method Var. Con. Elim. NNZ/Con. Lin. NZ/Con. Hess. NNZ
DIST 30368 30068 0 3.88 1.69 48032
LD1 30300 30000 68 3.89 1.69 48000
ECD2 29400 29100 968 3.62 1.68 47400
LD2 29068 28768 1300 3.64 1.67 47400
D2 19468 19168 10900 3.97 2.00 47100
GR 9932 9632 20436 4.92 1.00 47700
LM 9900 9600 20468 4.93 1.00 47700
MB 870 869 0 3.09 1.39 1869
LD1 780 779 90 2.97 1.34 1451
ECD2 777 776 93 2.95 1.34 1366
LD2 592 591 278 3.24 1.28 1285
D2 491 490 379 3.38 1.34 1145
GR 300 299 570 5.24 0.80 1609
LM 167 166 703 9.23 0.02 1524
OPF 61349 87120 0 3.06 2.27 75323
LD1 58969 89500 2380 2.90 2.16 72981
ECD2 55891 92578 5458 2.81 2.09 70249
LD2 55567 92902 5782 2.80 2.08 70113
D2 50650 97819 10699 2.65 1.64 70063
GR 27152 121317 34197 3.20 2.41 149351
LM 10396 138073 50953 3.93 0.24 80021
PIPE 12293 12221 0 3.00 1.70 33506
LD1 10372 10411 1921 2.56 1.72 11155
ECD2 7281 7731 5012 2.67 1.68 9414
LD2 5252 6281 7041 2.74 1.73 9414
D2 4660 5689 7633 2.81 1.80 7708
GR 5492 6834 6801 3.33 0.97 16627
LM 1124 2898 11169 8.26 0.92 25393
Table 4: Number of aggregations compared with bounds for the linear-matching strategy
Model Method Lower bound Number of Eliminations Upper bound
DIST LM 1868 20468 30068
MB LM 703 703 703
OPF LM 49794 50953 51488
PIPE LM 11169 11169 11197
Figure 5: Percent of variables eliminated by each method for each model. Methods are in the same order as presented in Table 3
We first note that the different aggregation methods eliminate significantly different numbers of variables for each model. Simply eliminating fixed variables can remove up to 10% of the variables in a model (e.g. moving bed and pipeline problems), while using an approximate-maximum aggregation strategy eliminates up to 90% of a model’s variables. The linear-degree-2 strategy eliminates anywhere from 4% of a model’s variables (in the distillation model) to 57% of a model’s variables (in the pipeline model). The percent of variables eliminated by each model-method combination is illustrated in Figure 5. For a given strategy, the wide variation in fraction of variables eliminated among different models suggests that we may expect significantly different runtime and reliability results when applying the same aggregation strategy to different models.
Figure 6: Incidence matrix for the moving bed optimization problem before variable aggregation and after variable aggregation using linear matching
The upper and lower bounds on number of aggregations performed by the linear-matching strategy according to Theorem 3.2 are shown in Table 4. In most cases, the linear-maximum aggregation strategy comes close to the upper bound on the number of eliminated variables, which is computed by performing a maximum matching on the linear-incidence graph. Only for the distillation column model is there a gap larger than 1%. For the distillation model, the block triangular form induced by the linear maximum matching is not highly decomposable, as indicated by the relatively low lower bound. However, the greedy procedure for recovering a lower triangular incidence matrix recovers an aggregation set with cardinality 68% of our upper bound. In the moving bed model, by contrast, the linear maximum matching induces a perfectly decomposable block triangular form. The OPF and pipeline models have maximum matchings that lead to highly decomposable block triangular forms. For the pipeline problem, the greedy procedure recovers only one variable and constraint per block in the block triangularization, yielding a number of eliminations equal to our lower bound. This is because, for the particular maximum matching computed, all diagonal blocks in the block triangularization of size either $1\times 1$ or $2\times 2$. In the latter case, only one variable can possibly be recovered, so the greedy procedure for recovering a lower triangular matrix cannot improve upon the lower bound.
Increased density of Jacobian and Hessian matrices in the aggregated problem is a concern that has been expressed by Gill et al. gill1981 and Achterberg et al. achterberg2020 . For this reason, most of our aggregation strategies are chosen with the intention of not increasing density in the remaining constraints. Indeed, density stays approximately constant after applying linear-degree-1, equal-coefficient-degree-2, linear-degree-2, or degree-2 strategies. However, the approximate-maximum aggregation strategies do not consider any limits on the degree of eliminated variables and constraints nodes, so they may lead to large increases in problem density. In the moving bed and pipeline problems, this strategy increases the number of variables incident on each constraint by factors of almost three, while for the other problems, the increase in density is more modest. Figure 6 shows the increase in density of the incidence matrix due to variable aggregation using linear matching. Additionally, the size of the incidence matrix decreases due to variable aggregation as variable-constraint pairs are eliminated from the problem. Although the density of the incidence matrix increases, it is still a relatively sparse matrix.
Figure 7: Number of linear and non-linear non-zeros per constraint after variable aggregation by each method for each model. Methods are in the same order as presented in Table 3
In addition to increased density of remaining constraints, aggressive variable aggregation may result in optimization problems that are “more nonlinear.” We quantify the increase in nonlinearity by measuring the average number of variables appearing linearly and nonlinearly in each constraint. Figure 7 demonstrates that for all the optimization problems, the number of linear variables per constraint decreases significantly when the linear-matching aggregation algorithm is used, even while the total number of nonzeros per constraint increases. On the other hand, the structure-preserving strategies approximately preserve the total non-zeros per constraint for all test problems.
5.2 Runtime results
Figure 8: Solve time per iteration spent in each category by each method for each model. Methods are in the same order as presented in Table 5. The time is normalized by the fastest solve time so that the fastest aggregation strategy has a normalized solve time per iteration of $1.0$.
Variable aggregation can also have a significant effect on an optimization problem’s solve time. We note that the direction of impact of variable aggregation on solve time is not immediately obvious. Aggregating variables leads to a smaller, denser optimization problem with potentially more inequality constraints (if many bounded variables are eliminated) and potentially more nonlinear constraints (if many variables are replaced with nonlinear expressions). The dominant computational costs in an interior point method are function and derivative evaluations and KKT matrix factorization. These are less expensive for a smaller problem, but KKT matrix factorization is more expensive for a denser problem, and Hessian evaluation is more expensive for a denser problem with more nonlinear constraints. For these reasons, it is important to study the breakdown of runtime to understand which factors are contributing to speedup or slow-down caused by variable aggregation. Finally, an interior point method may take significantly different numbers of iterations with full-space and aggregated optimization problems. While it is difficult to predict which formulation will converge in fewer iterations, it is important to observe trends that occur for a specific problem or problem class.
In this section we consider each problem’s solve time with IPOPT ipopt when using each of the proposed aggregation strategies. The times spent aggregating variables, solve times, iteration counts, and solve time breakdowns are given in Table 5.
In Table 5, $t_{\rm build}$ is the time to build the Pyomo optimization model, $t_{\rm elim}$ is the time to compute the sets of variables and constraints to aggregate and perform the aggregation in-place, $t_{\rm init}$ is the time to initialize data structures required for the optimization solve, $t_{\rm solve}$ is time spent by IPOPT solving the optimization problem, and “Iter.” is the number of iterations reported by IPOPT. “Func.”, “Jac.”, and “Hess.” are percentages of solve time spent in callbacks used by IPOPT for function, Jacobian, and Hessian evaluations. “Other” is the percentage of solve time spent in IPOPT itself (rather than one of these callbacks), which we assume is dominated by KKT matrix factorization. [^3] This is difficult to obtain programmatically, but by inspecting IPOPT’s logs we see, in the unmodified distillation problem for example, that IPOPT spends 70% of its time (i.e., 70% of the “Other” category) in a routine called “LinearSystemFactorization.” These percentages may not add up to 100 due to rounding.
Table 5: Runtime statistics of models after applying each aggregation method
Model Method Total times (s) Iter. Solve time breakdown (%) Objective
$t_{\rm build}$ $t_{\rm elim}$ $t_{\rm init}$ $t_{\rm solve}$ Func. Jac. Hess. Other
DIST 0.8 2.7 3.4 19 3 3 7 87 35.4
LD1 31.8 2.3 3.0 19 3 3 8 86 35.4
ECD2 38.3 2.4 1.6 17 5 5 12 78 35.4
LD2 38.0 2.4 1.9 18 5 5 14 76 35.4
D2 44.4 2.1 1.5 13 4 3 21 72 35.4
GR 17.7 2.1 1.5 13 5 3 24 69 35.4
LM 29.3 2.1 1.5 13 5 3 25 68 35.4
MB 4.1 0.5 0.1 9 10 6 19 65 0.00
LD1 2.7 0.5 0.1 9 10 7 19 63 0.00
ECD2 2.7 0.5 0.1 9 12 6 20 62 0.00
LD2 3.0 0.5 0.1 9 11 6 21 62 0.00
D2 3.3 0.4 0.1 10 12 6 27 56 0.00
GR 1.0 0.5 0.1 10 12 6 31 51 0.00
LM 1.5 0.5 0.1 10 11 5 49 35 0.00
OPF 2.8 4.3 9.8 61 4 4 9 82 1.39E+06
LD1 65.4 4.1 9.8 61 4 4 9 82 1.39E+06
ECD2 69.0 4.5 12.3 79 4 4 10 82 1.39E+06
LD2 76.7 4.5 44.9 198 3 3 7 87 1.39E+06
D2 84.5 4.6 42.5 200 3 4 10 82 1.39E+06
GR 30.7 5.0 16.8 75 4 5 10 82 1.39E+06
LM 62.5 8.1 27.4 73 9 11 26 54 1.39E+06
PIPE 6.0 1.4 6.8 49 1 1 4 95 6.27E+03
LD1 21.5 1.1 1.3 42 2 2 5 90 6.27E+03
ECD2 24.8 1.2 1.6 43 2 2 5 92 6.27E+03
LD2 25.2 1.4 1.1 43 2 2 7 89 6.27E+03
D2 26.0 1.1 1.2 43 3 2 8 88 6.27E+03
GR 7.0 1.3 1.4 30 3 2 9 87 6.27E+03
LM 29.9 1.8 3.5 40 2 4 75 19 6.27E+03
Figure 8 shows the normalized solve time per iteration in each category for every method and every optimization model. In two of the four problems considered (distillation and pipeline), variable aggregation significantly reduces solve time. However, even for these two problems, solve time does not uniformly decrease as more variables are eliminated. For example, the pipeline model takes 3.6 s to solve with the linear-matching strategy, but only 1.3 s to solve with the linear-degree-2, degree-2, or greedy strategies. For the OPF model, variable aggregation generally increases solve time, either due to larger iteration counts or an increase in the number of inequality constraints. For the moving bed model, variable aggregation does not have a measurable impact on solve time.
While variable aggregation can cause a significant reduction in solve time, it can also lead to expensive Hessian evaluations, introducing a new computational bottleneck. For all four models, the linear-matching strategy leads to a significantly larger fraction of solve time spent in Hessian evaluation, with a general trend that Hessian evaluation time (as a fraction of solve time) increases as more variables are eliminated (Figure 8). When solving the pipeline model after applying the linear-matching aggregation strategy, Hessian evaluation is responsible for 75% of total solve time and is the clear computational bottleneck in this solve. We note that our implementation takes advantage of the “defined variable” data structure in the ASL to not duplicate expressions when substituting an expression for a variable many times. Increases in Hessian evaluation times are due to the changing structure of expression graphs, rather than an increase in total size of expression graphs.
In contrast to Hessian evaluation time, the time spent in IPOPT (“Other” in Table 5) decreases as more variables are eliminated. This suggests that KKT matrix factorization benefits from the (generally) smaller systems in the aggregated problems and is not significantly slowed down by the accompanying increase in density.
5.3 Convergence reliability
As optimization problems before and after aggregation have different constraints, objectives, and derivative matrices, optimization algorithms may take significantly different paths through variable-space to arrive at solutions of the two problems. In some cases, one formulation may fail to converge within a specified iteration limit. We say that the formulation that converges more often is more reliable. As suggested by Parker et al. parker2022 , improving convergence reliability may be a reason to aggregate variables in nonlinear optimization problems.
In this section, we compare convergence reliability of each of our proposed elimination methods for distillation, moving bed, and pipeline test problems. We do not perform a parameter sweep for the OPF test problem as Egret knueven2019 does not construct these problems with easily mutable parameters. Here, reliability is measured for each test problem by performing a parameter sweep over two input parameters and counting the instances that are able to solve within 3,000 iterations with the IPOPT nonlinear solver.
Convergence results of a parameter sweep over relative volatility and the feed mole fraction of heavy component for the distillation optimal control problem are shown, for each method, in Figure 9. Results of a parameter sweep varying inlet solid temperature and inlet solid flow rate in the moving bed model are shown in Figure 10 and results of a sweep varying gas temperature and supply pressure in the pipeline dynamic optimization problem are shown in Figure 11. A summary showing the percent of problems solved by each method, for each model and in total, is shown in Table 6.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 9: Convergence of the distillation column optimal control problem after performing different aggregation strategies
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Convergence of the moving bed reactor optimization problem for different aggregation strategies
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 11: Convergence of the gas pipeline optimization problem for different aggregation strategies
Table 6: Percent of instances converged for methods applied to each model
Method Distillation Moving bed Pipeline Avg % convergence
% conv $t_{solve}^{avg}$ conv $t_{solve}^{avg}$ conv $t_{solve}^{avg}$
85 73.2 73 0.7 64 49.0 74
LD1 86 74.1 74 0.9 82 10.4 80
ECD2 88 45.9 74 0.7 82 6.9 81
LD2 90 56.1 73 0.9 82 6.5 82
D2 100 91.7 79 0.7 82 6.1 87
GR 98 78.5 87 0.8 82 6.3 89
LM 98 88.1 83 0.7 81 15.9 88
Our first observation is that variable aggregation generally improves convergence. No aggregation method leads to fewer instances converged, for any test problem, than the original model (represented by “–” in the “Model” column of Table 6). For each model, the discrepancy between instances converged with the original model and with the best aggregation method is significant: With the distillation model and the degree-2 strategy, 18 additional instances (15% more) converge, with the moving bed model and the greedy strategy, 17 additional instances (14% more) converge, and with the pipeline model and several different aggregation strategies, 22 additional instances (18% more) converge. Overall, the best strategies are degree-2 (Algorithm 4), greedy (Algorithm 1), and linear-matching (Algorithm 2), which converge between 13% and 15% more instances of these three test problems than the original model.
Here, methods that aggregate more variables generally converge more often, but the trend is not monotonic as the greedy strategy outperforms the linear-matching strategy despite aggregating significantly fewer variables for most models. Despite some methods performing significantly better than others on average, the improvement is not uniform for each problem instance. That is, there are some instances that fail with a “better” method despite succeeding with a “worse” method. We note that the reason for success or failure of an interior point method with variables aggregated or disaggregated is not easy to determine and believe that investigating the contribution of variable aggregation to enlarging or shrinking the basin of attraction of the interior point solution is an open and interesting area for future research. We believe that some of the improvement is due to primal iterates that stay closer to the feasible set of the original model and are more likely to have well-conditioned constraint Jacobians and KKT matrices than points along the infeasible path taken by an interior point method when solving the original model. We note that this explanation is similar to the motivation for IPOPT’s feasibility restoration phase (see Section 3.1 of ipopt ).
Refer to captionRefer to captionRefer to caption
Figure 12: Virtual best for all aggregation strategies for distillation, gas pipeline and moving bed reactor model
In algorithm design, the “virtual best” performance is defined as the performance of a theoretically perfect algorithm for variable aggregation. In this case, the virtual best algorithm picks the aggregation strategy from among the ones mentioned in Section 3 to maximize the convergence across the parameter sweep for each model. Figure 12 represents the virtual best for distillation, gas pipeline and the moving bed reactor models. For distillation, the degree-2 aggregation strategy is the virtual best since it achieves the same convergence. For the pipeline problem, all aggregation strategies except linear matching are virtual best. However, for the moving bed reactor, the virtual best achieves $90\%$ convergence while greedy—the best performing aggregation strategy—only achieves $87\%$ convergence. This indicates that there is further scope for improvement in the variable aggregation strategies to improve convergence reliability.
6 Conclusion
Different variable aggregation strategies are compared across four test problems in terms of structure, solve time and convergence reliability. The aggregation strategies are grouped into two main categories:
  • Structure-preserving strategies that aggregate variables without increasing the density of the remaining constraints
  • Approximate maximum aggregation strategies that maximize the variables aggregated but may lead to significant increase in the number of non-zeros in the remaining constraints
Different aggregation strategies aggregate significantly different number of variables across all test problems. The approximate maximum aggregation strategies, specifically the linear matching-based heuristic, eliminate a significantly higher percentage $(70-90\%)$ of variables compared to the structure-preserving strategies which eliminate $<60\%$ variables. However, it is important to note that the number of non-zeros per constraint stays approximately the same when structure-preserving aggregation strategies are used, while the approximate maximum aggregation strategies lead to an increase in the total number of non-zeros per constraint and also the increase the nonlinear non-zeros making the problem more nonlinear.
Variable aggregation can lead to a decrease in solve time as it decreases the factorization time of the KKT matrix. However, there are no guarantees that variable aggregation will decrease solve time as demonstrated in the OPF case where solve time increases after aggregating more variables due to an increase in inequality constraints introduced after elimination of bounded variables. Our work also suggests that Hessian evaluation can become a bottleneck if more variables are aggregated without preserving the structure of the problem, and hence, structure-preserving aggregation strategies are valuable.
The convergence reliability experiments indicate that variable aggregation leads to more reliable problem formulations. For the distillation example, degree-2 aggregation strategy leads to $100\%$ convergence over the parameter sweep. One reason for improved convergence may be that the aggregated problems follow a “more feasible” path with respect to the constraints of the original problem, leading to a better-conditioned KKT matrix. It is also observed that degree-2, greedy, and linear-matching strategies lead to more reliable convergence than the linearity-preserving, structure-preserving strategies. This result reveals an important trade-off: While aggregating many variables may introduce a new bottleneck due to expensive Hessian evaluation, it may lead to significantly more reliable convergence for some problems. For this reason, we recommend the degree-2 aggregation method, which strikes a balance between preserving structure and aggregating many variables, as a general pre-solve reduction.
While we observe that aggregation generally leads to more reliable convergence, we note that (a) the improvement is not uniform as more variables are eliminated and (b) there may be other problems for which aggregation causes worse convergence reliability. For instance, Albersmeyer and Diehl albersmeyer2010lifted suggest “lifting” a problem into a higher-dimensional space to improve iteration counts. Explaining the effects of problem reformulations on algorithms’ convergence properties is a significant challenge for nonlinear optimization, and this work will serve as a basis for further comparison and explanation of convergence behavior.
Based on our results, we believe that approximate-maximum and incidence-preserving aggregation methods should be implemented as pre-solve options in nonlinear modeling environments and solvers. However, further explanation and confirmation of our results is necessary to determine whether these aggregation methods are sufficiently general to be applied by default. We believe that the following efforts should be undertaken to explain and confirm the trends we observe:
  1. 1.
    Theoretical analysis of interactions among solvers’ globalization methods and different aggregation strategies.
  2. 2.
    Development of a test set of challenging, parameterized nonlinear optimization problems for which convergence of state-of-the-art solvers is not 100% reliable.
With these efforts, the methods we present may be further analyzed, theoretically and experimentally, and more informed development of nonlinear optimization software may proceed.
Appendix A Block triangularization
The block triangularization algorithm takes an undirected bipartite graph $G=(A,B,E)$ and a perfect matching thereof $\mathcal{M}$ as inputs and returns an ordered partition of $\mathcal{M}$, $\mathcal{B}$, that defines the irreducible block lower triangular form. Algorithms for permuting a sparse matrix or bipartite graph to irreducible block triangular form are well described by Duff and Reid duff1978implementation ; duff1978algorithm and Pothen and Fan pothen1990 . These descriptions define the algorithm as returning partitions of node sets $A$ and $B$, or of rows and columns of the sparse matrix. However, to facilitate the presentation of the matching-based aggregation algorithm, Algorithm 2, we define the block triangularization algorithm to return a partition of edges. For completeness, we provide a full description of the block triangularization algorithm in this appendix. The algorithm is defined using operations on a directed graph.
A directed graph $G=(V,E)$ is a graph in which edges are ordered pairs of nodes. If $E$ contains an edge $(u,v)$, $v$ is said to be out-adjacent to $u$, while $u$ is in-adjacent to $v$. In a directed graph, a path is a sequence of nodes $\{u_{1},\dots,u_{n}\}$ where, for every adjacent pair in the sequence $(u_{i},u_{i+1})$, $u_{i+1}$ is out-adjacent to $u_{i}$. A strongly connected component $C$ is a subset of nodes in a graph where, for every pair of nodes $u,v\in C$, there exists a path from $u$ to $v$. The set of all strongly connected components $\mathcal{C}$ partitions the nodes in a directed graph. Strongly connected components of a directed graph may be computed in $\mathcal{O}(n_{v}+n_{e})$ time, where $n_{v}$ is the number of vertices and $n_{e}$ is the number of edges, by Tarjan’s algorithm tarjan1972 . A cycle is a path that starts and ends with the same node, and a directed acyclic graph (DAG) is a directed graph that has no cycles. The strongly connected components of a directed graph form a DAG that can be constructed by the compress subroutine described in Algorithm 9. A topological order is a (nonunique) permutation of the nodes in a DAG such that, if $(u,v)$ is an edge in the DAG, $u$ comes before $v$ in the topological order. A topological order may be computed in $\mathcal{O}(n_{e}+n_{v})$ time manber1989 .
The block triangularization algorithm is given by Algorithm 7. It consists of five major steps:
  1. 1.
    Project the bipartite graph into a directed graph defined over the bipartite set $A$, using the matching $\mathcal{M}$ to determine edge orientation.
  2. 2.
    Partition the nodes $A$ into strongly connected components of the new directed graph.
  3. 3.
    Compress the strongly connected components into a DAG. Each node of the DAG is a subset of nodes in the directed graph (a subset of $A$).
  4. 4.
    Compute a topological order of the DAG.
  5. 5.
    For each subset of $A$ in the topological order, convert each node $a$ to an edge $(a,b)$ by finding the matched vertex in $\mathcal{M}$.
Descriptions of the projection and compression subroutines are given in Algorithms 8 and 9, and Table 7 summarizes all subroutines used by the block triangularization algorithm. By construction, the returned set of sets of edges, $\mathcal{T}$ in Algorithm 7, is a partition of the perfect matching $\mathcal{M}$.
Table 7: Subroutines used by block_triangularize algorithm
Subroutine Inputs Outputs Time complexity Description
adjacent_to Graph $G$, node $u$ Set of nodes $\mathcal{O}(\delta(u))$ Set of nodes in $G$ adjacent to $u$
matched_with Matching $\mathcal{M}$, node $u$ Node $\mathcal{O}(1)$ Node matched with $u$ in $\mathcal{M}$
subset_containing Set of disjoint sets, node $u$ Set $\mathcal{O}(1)$ Return set containing $u$
project Bipartite graph, matching Directed graph $\mathcal{O}(n_{e})$ Directed graph defined on one bipartite set
strongly_connected_comps Directed graph Set of sets of nodes $\mathcal{O}(n_{v}+n_{e})$ Return the strongly connected components of the graph
compress Graph, set of sets of nodes Graph $\mathcal{O}(n_{v}+n_{e})$ Each specified subset of nodes in the original graph is a single node in the returned graph
topological_sort Directed acyclic graph Ordered set of nodes $\mathcal{O}(n_{v}+n_{e})$ Permute the nodes into topological order
Algorithm 7 : block_triangularize
1:Inputs: Bipartite graph $G=(A,B,E)$, perfect matching $\mathcal{M}$
2:$G_{d}=\text{\tt project}(G,\mathcal{M})$triangleright The nodes of $G_{d}$ are $A$
3:$\mathcal{C}=\text{\tt strongly\_connected\_comps}(G_{d})$triangleright $\mathcal{C}$ partitions $A$
4:$D=\text{\tt compress}(G_{d},\mathcal{C})$triangleright The nodes of $D$ are $\mathcal{C}$
5:$\mathcal{S}=\text{\tt topological\_sort}(D)$
6:$\mathcal{T}=\{\}$
7:for $C\text{~{}\bf in }\mathcal{S}$ do $\triangleright$ $C$ is a subset of $A$
8:     $T=\{\}$triangleright $T$ is a subset of $\mathcal{M}$
9:     for a in $\text{~{}\bf in }C$ do
10:         $b=\text{\tt matched\_with}(\mathcal{M},a)$
11:         $T\leftarrow T\cup(a,b)$
12:     end for
13:     $\mathcal{T}\leftarrow\mathcal{T}\cup T$
14:end for
15:Return: $\mathcal{T}$
Algorithm 8 : project
1:Inputs: Bipartite graph $G=(A,B,E)$, matching $\mathcal{M}$
2:$E_{d}=\{\}$
3:for $a\text{~{}\bf in }A$ do
4:     $b={\tt matched\_with}(\mathcal{M},a)$
5:     for $\bar{a}\text{~{}\bf in }\text{\tt adjacent\_to}(b)$ do
6:         $E_{d}\leftarrow E_{d}\cup(\bar{a},a)$
7:     end for
8:end for
9:Return: $G_{d}=(A,E_{d})$
Algorithm 9 : compress
1:Inputs: Graph $G=(V,E)$, partition $\mathcal{C}$
2:$V_{\mathcal{C}}=\mathcal{C}$
3:$E_{\mathcal{C}}=\{\}$
4:for $C\text{~{}\bf in }\mathcal{C}$ do
5:     for $u\text{~{}\bf in }C$ do
6:         for $v\text{~{}\bf in }\text{\tt adjacent\_to}(u)$ do
7:              $\bar{C}=\text{\tt subset\_containing}(\mathcal{C},v)$
8:              if $C\neq\bar{C}$ then
9:                  $E_{\mathcal{C}}\leftarrow E_{\mathcal{C}}\cup(C,\bar{C})$
10:              end if
11:         end for
12:     end for
13:end for
14:Return: $G_{\mathcal{C}}=(V_{\mathcal{C}},E_{\mathcal{C}})$